TC(8) Linux TC(8) NAME choke−chooseandkeepscheduler SYNOPSIS tcqdisc...chokelimitpacketsminpacketsmaxpacketsavpktbytesburstpackets[ecn][ bandwidthrate]probabilitychance DESCRIPTION CHOKe(CHOoseandKeepforresponsiveflows,CHOoseandKillforunresponsiveflows)isa classlessqdiscdesignedtobothidentifyandpenalizeflowsthatmonopolizethequeue.CHOKeis avariationofRED,andtheconfigurationissimilartoRED. ALGORITHM Oncethequeuehitsacertainaveragelength,arandompacketisdrawnfromthequeue.Ifboth theto-be-queuedandthedrawnpacketbelongtothesameflow,bothpacketsaredropped.Oth-erwise,ifthequeuelengthisstillbelowthemaximumlength,thenewpackethasaconfigurable chanceofbeingmarked(whichmaymeandropped).Ifthequeuelengthexceedsmax,thenew packetwillalwaysbemarked(ordropped).Ifthequeuelengthexceedslimit,thenewpacketis alwaysdropped. ThemarkingprobabilitycomputationisthesameasusedbytheREDqdisc. PARAMETERS TheparametersarethesameasforRED,exceptthatREDusesbyteswhereaschokecountspackets.Seetc-red(8)foradescription. SOURCE o R.Pan,B.Prabhakar,andK.Psounis,"CHOKe,AStatelessActiveQueueManagement SchemeforApproximatingFairBandwidthAllocation",IEEEINFOCOM,2000. o A.Tang,J.Wang,S.Low,"UnderstandingCHOKe:ThroughputandSpatialCharacteristics",IEEE/ACMTransactionsonNetworking,2004 SEEALSO tc(8),tc-red(8) AUTHOR schedchokewascontributedbyStephenHemminger. iproute2 August2011 1